Grameen Phone (GP) is planning to finance a project using the following weights: 35% long term debt, 15% preferred stock and 50% common stock equity. The firm’s tax rate is 40%

I. The firm can sell for Tk. 970 and 12 years bond, Tk. 1000 par-value bond paying at a 13% coupon rate. A floatation cost of 2% of par value is required.

II. Eight percent (annual dividend) preferred stock having a par value of Tk. 100 can be sold for Tk. 88. An additional fee of Tk. 5 per share must be paid to the underwriters.

III. The firm’s common stock currently selling for Tk. 68 per share. The dividend expected to be paid at the end of the year is Tk. 5 per share. Dividend have been growing at an annual rate of 8%, firm must also pay Tk. 3 per share in flotation cost.

a. Calculate the WACC of the project and the significance of WACC for project investment decision.

b. Differentiate between independent and mutually exclusive projects.


1
Expert's answer
2020-12-24T17:20:41-0500

a. WACC find the formula:

"WACC=Re\frac{E}{V}+Rd(1-t)\frac{D}{V}"

where:

re – return on equity of the organization;

rd — return on borrowed capital of the organization;

E/V, D/V – share of equity and borrowed capital in the capital structure of the company.

The amount of equity and debt capital forms the company's capital (V=E+D);

t is the interest rate of income tax.

find interest rates on stocks and bonds:

I. "rd=\frac{\frac{N-P}{n}+C}{\frac{N+P}{2}}=\frac{\frac{1020-970}{12}+132.6}{\frac{1020+970}{2}}=\frac{136.77}{995}=13.75"

II.

"D=105\times0.08=8.4"


"P=\frac{D}{R}"

"88=\frac{8.4}{R}"

R=9.55

III. according to Gordon's model

"P=\frac{D}{R-g}"

P=68+3=71

"71=\frac{5}{R-0.08}"

R=15.04


"WACC=9.55\times0.15+15.04\times0.5+13.75\times(1-0.4)\times0.35=1.43+7.52+2.89=11.84"

b. A Project whose cash flows have no impact on the acceptance or rejection of other projects is termed as Independent Project (not mutually exclusive).

A set of projects from which at most one will be accepted is termed as Mutually Exclusive Projects.